OpenStudy (anonymous):

45% percent of what number is 4.5

OpenStudy (anonymous):

i think its 2.025

OpenStudy (anonymous):

what are the choices

OpenStudy (paki):

45/100*X= 4.5.... solve for "X"....

OpenStudy (jadeishere):

i think it's 10...

OpenStudy (jadeishere):

it's pretty simple,. actually, 4.5 out of 10 is 45

OpenStudy (texaschic101):

45% of what number is 4.5....write it like you read it.. of means multiply is means equals turn 45% into a decimal which is 0.45 0.45x = 4.5 -- divide by 0.45 x = 4.5/0.45 x = 10